OpenStudy (anonymous):

A plane has 240 gallons of fuel. It averages 3 3/8 miles per gallon. How far can the plane go?

OpenStudy (anonymous):

27*30 = 810 miles

OpenStudy (anonymous):

240 x 3.375 = 810 miles

OpenStudy (anonymous):

\[3+(3/8) = ((3\times 8)+3)/8 = 27/8\]\[240\times 27/8=(240\times 27)\div 8\]\[=6480\div 8\]\[=810\]

OpenStudy (anonymous):

\[240*\left(3+\frac{3}{8}\right)\text{= }810 \text{ miles}\]